OpenStudy (anonymous):

A number is increased by 70%, or 210. What is the original number?

jimthompson5910 (jim_thompson5910):

70% of original number = 210 70% of x = 210 (70/100)*x = 210 0.7x = 210 I'll let you finish

OpenStudy (anonymous):

140?

jimthompson5910 (jim_thompson5910):

0.7x = 210 x = 210/0.7 x = ???

OpenStudy (anonymous):

300.

jimthompson5910 (jim_thompson5910):

good